A claim is made that the proportion of children who play a sport is less than 0.5. A sample is

taken of 500 children which finds that 40%

of them play sports.State your hypothesis and test

the claim at .1 significance.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p = 0.5

Ha : p < 0.5

\hat p = 0.40

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.40 - 0.5 / [\sqrt(0.5 * 0.5) / 500]

= -4.472

P-value = 0.0000

\alpha = 0.01

P-value < \alpha

Reject the null hypothesis .

There is sufficient evidence to suggest that  proportion of children who play a sport is less than 0.5.